If a student takes 5 quizzes and scores 79, 86, 74, 61 and 94, what does she need to score to have a 80 average?
Answer by stanbon(75887)

If a student takes 5 quizzes and scores 79, 86, 74, 61 and 94, what does she need to score to have a 80 average?
--------------
Let the 6th score be "x"::
Equation:
(79+86+74+61+94+x)/6 = 80
--------
394 + x = 6*80
----
x = 86 (score needed on the 6th test)
=============
Cheers,
Stan H.